# Towards Active Synthetic Data Generation for Finetuning Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable abilities in a wide variety of reasoning and factual knowledge tasks (Achiam et al., 2023; Bubeck et al., 2023; Katz et al., 2024), but their large size makes inference expensive. With the advent of agentic systems that interact with the external world, LLMs are poised to become even more ubiquitous in science, technology, and society, but the tremendous inference cost presents a challenge for realizing the full potential of these agents.
One way to quell the computational expense associated with LLM inference is to use small language models (SLMs). With orders of magnitude fewer parameters, SLMs are faster, cheaper, and easier to finetune for specialised skills like tool use, making them natural specialists using proprietary data or within agentic systems (Belcak et al., 2025).
Training language models typically involves three stages: pre-training on large general-purpose corpora, supervised finetuning (SFT), and reinforcement learning from human feedback (RLHF) or from verifiable rewards (RLVR) (Ouyang et al., 2022). SFT, the focus of this work, is critical for
Figure 1: Overview of iterative synthetic data generation (Algorithm 1). The student model guides synthetic data generation by prioritizing which data are used as an example for the teacher model to generate a new synthetic data point (Section 4.2). The student finetunes on teacher generated synthetic data.
@@
However, real-world data for SFT can be hard to obtain, or may lack desirable properties such as chain-of-thought reasoning (Wei et al., 2022). Consequently, a typical strategy involves synthesizing a corpus of question and answer pairs from a larger, more capable model (Mitra et al., 2024; Liu et al., 2024a). This process usually begins with a small seed dataset and leverages a teacher LLM to produce supplementary synthetic samples, then finetunes the student SLM on the resulting sequences in aggregate.
Still, evidence suggests that generating a large static synthetic dataset is frequently wasteful, as it can often be drastically pruned with little to no degradation in trained model capabilities  (Chen et al., 2023; Zhou et al., 2024). As such, this paper explores an iterative, targeted approach to synthetic data generation that is student-aware and improves data efficiency—achieving stronger performance under a fixed data generation budget than naive, static generation—thereby yielding a superior performance–training-set-size Pareto frontier (see Section 2 for a formal definition).
@@
  • We provide a benchmark study for iterative synthetic data generation rooted in prior work on active learning and data selection. We carefully compare to static dataset generation—identical to uniform random sampling of prompts—to show improvements in data efficiency.
  • We compare a range of methods for selecting samples for synthetic data generation, including those that favour uncertainty, diversity, or difficult/easy samples. We conclude that simple methods rooted in active learning, such as using the loss of the student’s own prediction (Settles and Craven, 2008) are the most data efficient. In contrast, expensive and popular methods based on using an LLM to judge the difficulty and quality of data, i.e. LLM-as-a-judge (Zheng et al., 2023; Jiang et al., 2023b), underperform simpler and more general active learning alternatives.
  • We show that synthetic data generation is to a certain extent “steerable”; the properties of the teacher generated synthetic data resemble those of the data selected by the student. For example, if the student selects “difficult” data—with a high loss—then the teacher also generates “difficult” data in aggregate.
@@ 3 Related Work
Synthetic datasets are known to be compressible—synthetic samples filtered to have high LLM-as-a-judge (Chen et al., 2023) values or low student loss (Li et al., 2024), for example, can obtain the same performance as finetuning on the entire unpruned corpus. To remedy this inefficiency, rather than generating a large static synthetic dataset and then filtering, we can instead carefully select the seed data used to generate the synthetic samples, and hopefully produce fewer semantically similar sequences. This has been shown by prioritizing incorrect samples when querying the teacher, which is more data efficient than finetuning on the original corpus $D_{0}$ (Lee et al., 2024). LLM-as-a-judge selection is also known to be more data efficient than directly finetuning on a public benchmark synthetic datasets (Jiang et al., 2023b). This study includes LLM-as-a-judge scoring due to its widespread use and prioritizing incorrectly answered student responses due to its simplicity.
@@ 3.1 Assigning a value to data
Our work makes use of ideas from active learning, which seeks to maximise data efficiency by iteratively identifying and prioritising informative samples for labelling (Settles, 2009; Settles and Craven, 2008). Classic strategies for active learning include model prediction disagreement (Freund et al., 1997; Houlsby et al., 2011), uncertainty (MacKay, 1992; Gal et al., 2017; Kirsch et al., 2019), and dataset summarization (Sener and Savarese, 2018; Mirzasoleiman et al., 2020; Coleman et al., 2019). Effective, contemporary methods trade-off between predictive uncertainty and sample diversity in a fashion that is commensurate with large neural networks (Ash et al., 2021; Saran et al., 2023). We consider language model-aligned variations of two popular methods for active learning: uncertainty sampling (Settles and Craven, 2008), a classic approach that favours predictive uncertainty, and BADGE, a more modern algorithm (Ash et al., 2019).
@@
Related methods aim to estimate the value of data to guide selection, typically using labelled dataset $(x,y)$. Data can be valued using Shapley values (Ghorbani and Zou, 2019), influence functions Koh and Liang (2017) or by matching training data to desirable evaluation datasets Just et al. (2023); Kessler et al. (2025); these methods have shown limited effectiveness for language modelling. LLMs have been used to score data points (Zheng et al., 2023) and for selecting question-answer samples for SFT (Liu et al., 2024b; Jiang et al., 2023b; Chen et al., 2023). Still, it has been shown that LLMs scores exhibit biases that hinder their effectiveness in this setting (Xiong et al., 2024; Dorner et al., 2025; Panickssery et al., 2024). Alternative approaches use training loss or gradient norms with respect to student parameters as an estimate of learning progress (Loshchilov and Hutter, 2015; Katharopoulos and Fleuret, 2018; Jiang et al., 2019; Li et al., 2024; Mindermann et al., 2022; Evans et al., 2024; Dai et al., 2025). However, this has shown limited data efficiency for language models (Kaddour et al., 2023). Reward models are commonly used to score and identify data points for SFT (Cao et al., 2023; Dubey et al., 2024). This work focuses on reward selection because of its popularity.
@@ 4.1 Selection Algorithms
A common method in the active learning literature is uncertainty sampling, which, for non-sequential classification models, prioritizes data for which the amount of probability mass on the most likely class predicted by the model is smallest (MacKay, 1992). In the sequential, Transformer-based setting, we can score a data point with the loss of the response tokens under the student $f_{{\bm{\theta}}}(\cdot)$ with parameters $\bm{\theta}$} as $\mathcal{L}(z_{i},{\bm{\theta}})$ (Settles and Craven, 2008). When the targets used to produce a loss are the model’s own generations, this score reflects an uncertainty in the produced sequence. Note that our setting gives us access to the ground-truth label associated with $x$ as well, and thus allows us to compute a true loss in a fashion commensurate with conventional model training (Loshchilov and Hutter, 2015). Interestingly, we find this to be less effective empirically than using the former, uncertainty-based approach.
@@
We can also leverage the reasoning ability of an LLM teacher model to score an SLM’s predictions. This strategy asks the LLM teacher to score the detail, quality and correctness of the student’s answer and reasoning with a value between $[1,10]$. In particular, we use pairwise LLM-as-a-judge scoring which has been shown to be most effective (Zheng et al., 2023). Two separate answers are given for the teacher to decide which it prefers by providing scores for both: $s_{i}^{t},s_{i}=\texttt{LLM}(\hat{y}^{t}_{i},\hat{y}_{i},x_{i})$ where $\hat{y}^{t}_{i}=\texttt{LLM}(x_{i})$ is teacher’s answer, $s_{i}^{t}$ is the score for the teachers answer and $\hat{y}_{i}$ the student answer. This is an expensive scoring method since it requires the teacher to produce an answer in addition to scoring.
@@ 4.2 Prompt-based synthetic data generation
Selected data points $\bar{x}_{i}\in\bar{D}_{t}$ are added to a synthetic data generation prompt for the LLM teacher model to generate a synthetic question $\hat{x}_{i}$ (Xu et al., 2024; Mitra et al., 2024; Jiang et al., 2023b; Lee et al., 2024). The teacher is then prompted to produce chain-of-thought reasoning and a final answer for $\hat{y}_{i}$. We generate a synthetic data point $\hat{z}_{i}=(\hat{x}_{i},\hat{y}_{i})$ using $\hat{x}=\texttt{LLM}(\bar{x}_{i})$ and $\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})$. So $\hat{D}_{t}=\text{Generate}(\bar{D}_{t})=\left\{\hat{x}_{i}=\texttt{LLM}(\bar{x}_{i}),\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})\right\}_{i=1}^{m}\text{, where}\,\bar{x}_{i}\sim\bar{D}_{t}$. For the further details on the prompts used for each dataset see Section D.2.
@@ 5 Experiments
This section empirically probes the data efficiency of iterative synthetic data generation against static data generation, and provides recommendations regarding which scoring and selection design choices improve efficiency. We find that prioritizing difficult data, measured as the student’s loss on its own generation, to be most data efficient. We further provide probing experiments that give an exploration for this improved data efficiency. We show that on average synthetic data inherits some properties from the samples used to generate them. If we choose difficult samples for the student—measured by a high loss or a low reward for example—the synthetic data is difficult as well, resulting in lower student accuracies on these generated samples than random selection. Both of these observations explain why selecting data prior to synthetic data generation results in synthetic data that has similar properties to our selected data.
At each iteration $t$ we use a given acquisition algorithm to select 11k samples $\bar{D}_{t}$ from $D_{t}$, before sending each to the teacher model to generate corresponding synthetic data $\hat{D_{t}}$. These data are appended to synthetic data from all previous iterations before reinitializing the student model and refitting its parameters.
@@ 5.1 Datasets
This section presents results on four distinct reasoning datasets in conjunction with four different models. GSM8k is a popular mathematics dataset comprised of school level maths problems (Cobbe et al., 2021), which we use with a Mistral-7B-Instruct-v0.3 student (Jiang et al., 2023a). Similarly, we include the more challenging Math1-3 dataset (Hendrycks et al., 2021), which is comprised of $5$ distinct levels of question difficulty—we use the easiest levels, $1$ to $3$, to finetune a Llama-3-8B-Instruct student (Dubey et al., 2024). We further experiment with the logical reasoning dataset ProntoQA (Saparov and He, 2023), composed of synthetically generated chain-of-thought style reasoning questions, with a Qwen1.5-7B-Chat student. Finally we consider the Game of 24 dataset, where a model is required to find the arithmetic operations to obtain $24$ given $4$ separate numbers. Here we use a Qwen2.5-7B-Instruct student (Qwen et al., 2025). More dataset details are provided in Section D.1.
For all datasets except for Game of 24 we use prompt-based synthetic data generation with a GPT-4o teacher; see Section D.2 for our prompts. For Game of 24 we use backward reasoning: if the answer is 13*8-10*8=24, for example, we can construct a new question by setting two integers to variables a*b-10*8=24 and solving to generate new questions (Jiang et al., 2024). We use a GPT-o3-mini teacher for backward reasoning, as it seems to produce better question-response pairs than GPT-4o (Section D.2.4).
@@ 5.2 Finetuning Setup
Figure 2: SFT performance on 11k data points for various datasets and SLMs. We compare the effect of synthetic answer generation and synthetic question and answer generation to using the seed dataset, $D_{0}$ for SFT. 0-shot SLM and teacher performances are included for reference. All datasets use a GPT-4o teacher, for Game of 24 we use a GPT-o3-mini teacher.
@@ 5.3 Algorithms
We consider a variety of selection algorithms to understand which are most sample efficient. Prior work has shown that prioritizing “hard” samples accelerates learning (Section 3.1), which we also find to be the case for iterative synthetic data generation (Section C.6). This approach prioritizes high-uncertainty data, measured as the model’s loss on greedily decoded student generations, which we denote as “loss (high)” throughout this section. We also consider a low-reward selection algorithm (“rwd (low)”), also using the student’s own predictions, which scores generations using an external model. We use a Skywork-Reward-Llama-3.1-8B-v0.2 reward model which is the highest scoring 88b model on RewardBench (Lambert et al., 2025) at the time of writing.
We use Lion (Jiang et al., 2023b) as a baseline, which compares the student and teacher answer LLM-as-a-judge scores to categorize a data point as either hard or easy. All seed data are assigned into either an easy or a hard set before sampling equally from both. For completeness, we also consider a baseline that only samples from the hard set, denoted as LLM-as-a-judge (hard) (Jazbec et al., 2024). We use the same prompts for LLM-as-a-judge scoring as Jiang et al. (2023b).
We further consider prioritizing data with incorrect student answers, $s_{i}=\bm{1}\{\hat{y}_{i}\neq y\}$ as a proxy for prioritizing hard samples (Lee et al., 2024). In a similar spirit to Lion, we can instead sample evenly from correct and incorrect pools to maintain diversity in the synthetic data used for generation (Liu et al., 2024a). Since correct and incorrect scoring requires a verifier and ground-truth answers, we do not compare them to other scoring methods which do not use label information and we place these supplementary results in Section C.1.
Figure 3: Student performance over successive synthetic data iterations with growing training sets (Algorithm 1). Each inset plot shows the proportion of data random sampling requires for the same performance as the best active scorers for synthetic data generation

### v2

@@ 1 Introduction
Despite the tremendous cost of inference, Large Language Models (LLMs) have risen to prominence as a result of their remarkable abilities across a wide array of reasoning and factual knowledge tasks (Achiam et al., 2023; Bubeck et al., 2023; Katz et al., 2024). As agentic systems capable of interacting with the external world emerge, these models are poised to become even more ubiquitous in science, technology, and society, but the tremendous inference cost presents a challenge for realizing the full potential of these agents.
One way to quell the computational expense associated with LLM inference is to use small language models (SLMs). With orders of magnitude fewer parameters, SLMs are faster, cheaper, and easier to finetune for specialised skills like tool use, making them natural specialists using proprietary data or within bespoke agentic systems (Belcak et al., 2025).
Training language models typically involves three stages: pre-training on large, general-purpose corpora, supervised finetuning (SFT), and reinforcement learning from human feedback (RLHF) or from verifiable rewards (RLVR) (Ouyang et al., 2022). SFT, the focus of this work, is critical for
Figure 1: Overview of iterative synthetic data generation (Algorithm 1). The student model guides synthetic data generation by prioritizing which data are used as an example for the teacher model to generate a new synthetic data point (Section 4.2). The student finetunes on teacher generated synthetic instruction data.
@@
In practice, real-world data for SFT can be hard to obtain, or may lack desirable properties such as chain-of-thought reasoning (Wei et al., 2022). Consequently, a typical strategy involves synthesizing a corpus of question and answer pairs from a larger, more capable model (Mitra et al., 2024; Liu et al., 2024a). This process usually begins with a small seed dataset, which a teacher LLM uses to produce supplementary synthetic samples before the student SLM is finetuned on the resulting sequences.
Still, evidence suggests that generating a large, static synthetic dataset is frequently wasteful, as it can often be drastically pruned with little to no degradation in trained model capabilities  (Chen et al., 2023; Zhou et al., 2024). As such, this paper explores an iterative, targeted approach to synthetic data generation that is student-aware and improves data efficiency—achieving stronger performance under a fixed data generation budget than naive, static generation—thereby yielding a superior performance–training-set-size Pareto frontier (see Section 2 for a formal definition).
@@
  • We provide a benchmark study for iterative synthetic data generation rooted in prior work on active learning and data selection, and compare to static synthetic dataset generation. We show improvements in data efficiency when comparing to generating a single large synthetic instruction dataset, which is a typical approach to student post-training (Mitra et al., 2024; Luo et al., 2025).
  • We compare a range of methods for selecting samples for seeding synthetic data generation, including those that favour uncertainty, diversity, or difficulty. We conclude that simple methods rooted in active learning, such as using the loss of the student’s own prediction are most data efficient. In contrast, expensive and contemporary methods that use an LLM to judge the difficulty and quality of data, i.e. LLM-as-a-judge (Zheng et al., 2023; Jiang et al., 2023b), surprisingly underperform when compared to simpler alternatives.
  • We show that synthetic data generation is, to a certain extent, steerable: properties of teacher-generated synthetic data resemble those used to seed the generation process. If the student selects challenging data—measured as samples that induce high student loss—the teacher generates data with correspondingly high loss on average. Given this relationship, and the fact that recent, specialised, LLM-based strategies often fall short, we argue that research in selection strategies is a fruitful and underexplored avenue for advancing the performance of small language models.
@@ 2 Preliminaries
$\mathcal{L}(z,{\bm{\theta}})=\nicefrac{{-1}}{{|y|}}\sum_{j=1}^{|y|}y_{j}\log f_{{\bm{\theta}}}(x,y_{<j}).$
The model $f_{{\bm{\theta}}}(\cdot)$ autoregressively generates the next token $\hat{y}_{j}=f_{{\bm{\theta}}}(x,\hat{y}_{<j})$ in the sequence.
@@ 3 Related Work
[…]
@@ 3.1 Assigning a Value to Data
Our work makes use of ideas from active learning, which seeks to maximise data efficiency by iteratively identifying and prioritising informative samples for labelling (Settles, 2009; Settles and Craven, 2008). Classic strategies for active learning include model prediction disagreement (Freund et al., 1997; Houlsby et al., 2011), uncertainty (MacKay, 1992; Gal et al., 2017; Kirsch et al., 2019), and dataset summarization (Sener and Savarese, 2018; Mirzasoleiman et al., 2020; Coleman et al., 2019). Effective, contemporary methods trade-off between predictive uncertainty and sample diversity in a fashion that is commensurate with large neural networks (Ash et al., 2021; Saran et al., 2023). We consider language model-aligned variations of two popular methods for active learning: uncertainty sampling (Settles and Craven, 2008), and BADGE, a more modern algorithm that trades-off between predictive uncertainty and the diversity of selected data (Ash et al., 2019).
@@
Related methods aim to estimate the value of data to guide selection, typically using a labelled dataset $(x,y)$. Data can be valued using Shapley values (Ghorbani and Zou, 2019), influence functions Koh and Liang (2017) or by matching training data to evaluation datasets Just et al. (2023); Kessler et al. (2025); these methods have shown limited effectiveness for language modelling. LLMs have been used to score data points (Zheng et al., 2023) and for selecting question-answer samples for SFT (Liu et al., 2024b; Jiang et al., 2023b; Chen et al., 2023). Still, it has been shown that LLM scores exhibit biases that hinder their effectiveness in this setting (Xiong et al., 2024; Dorner et al., 2025; Panickssery et al., 2024). Alternative approaches use training loss or gradient norms with respect to student parameters as an estimate of learning progress (Loshchilov and Hutter, 2015; Katharopoulos and Fleuret, 2018; Jiang et al., 2019; Li et al., 2024; Mindermann et al., 2022; Evans et al., 2024; Dai et al., 2025). However, this has shown limited data efficiency for language models (Kaddour et al., 2023). Reward models are commonly used to select data points for SFT (Cao et al., 2023; Dubey et al., 2024). This work focuses on reward selection because of its popularity.
@@ 4.1 Selection Algorithms
A common method in the active learning literature is uncertainty sampling, which, for non-sequential classification models, prioritizes data whose probability mass on the most likely class predicted by the model is smallest (MacKay, 1992). In the sequential, Transformer-based setting, we can score a data point with the loss of the response tokens under the student $f_{{\bm{\theta}}}(\cdot)$ with parameters $\bm{\theta}$} as $\mathcal{L}(z_{i},{\bm{\theta}})$ (Settles and Craven, 2008). When the targets used to produce a loss are the model’s own generations, this score reflects an uncertainty in the produced sequence. Note that our setting gives us access to the ground-truth label associated with $x$ as well, and thus allows us to compute a true loss in a fashion commensurate with conventional model training (Loshchilov and Hutter, 2015). Interestingly, we find this to be less effective empirically than using the former, uncertainty-based approach (Section C.6).
@@
We can also leverage the reasoning ability of an LLM teacher model to score an SLM’s predictions. This strategy asks the LLM teacher to score the detail, quality and correctness of the student’s answer and reasoning with a value between $[1,10]$. In particular, we use pairwise LLM-as-a-judge scoring which has been shown to be most effective (Zheng et al., 2023). Two separate answers are given for the teacher to decide which it prefers by providing scores for both: $s_{i}^{t},s_{i}=\texttt{LLM}(\hat{y}^{t}_{i},\hat{y}_{i},x_{i})$ where $\hat{y}^{t}_{i}=\texttt{LLM}(x_{i})$ is teacher’s answer, $s_{i}^{t}$ is the score for the teachers answer and $\hat{y}_{i}$ the student answer. This is expensive, as it requires the teacher to produce an answer in addition to scoring.
@@ 4.2 Prompt-based Synthetic Data Generation
Selected data points $\bar{x}_{i}\in\bar{D}_{t}$ are added to a synthetic data generation prompt for the LLM teacher model to generate a synthetic question $\hat{x}_{i}$ (Xu et al., 2024; Mitra et al., 2024; Jiang et al., 2023b; Lee et al., 2024). The teacher is then prompted to produce chain-of-thought reasoning and a final answer for $\hat{y}_{i}$. We generate a synthetic data point $\hat{z}_{i}=(\hat{x}_{i},\hat{y}_{i})$ using $\hat{x}=\texttt{LLM}(\bar{x}_{i})$ and $\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})$. So $\hat{D}_{t}=\text{Generate}(\bar{D}_{t})=\left\{\hat{x}_{i}=\texttt{LLM}(\bar{x}_{i}),\hat{y}_{i}=\texttt{LLM}(\hat{x}_{i})\right\}_{i=1}^{m}\text{, where}\,\bar{x}_{i}\sim\bar{D}_{t}$. For details on prompts used for each dataset see Section D.2.
@@ 5 Experiments
This section empirically probes the data efficiency of iterative synthetic data generation against static data generation, and provides recommendations for scoring and selection design choices for data efficiency. We find that prioritizing challenging data, as measured by the student’s loss on its own generations, to be at least as data efficient as teacher-based LLM scoring methods, and often more efficient.
LLM-based scoring can behave erratically, particularly for unusual tasks likely outside of the model’s training data distribution. Paired with the additional expense of using a large LLM to score data, more general approaches, like uncertainty sampling, appear to be more reliable and effective.
We further explore this improved data efficiency and show that on average synthetic data inherits some properties from samples used to generate them. If we select data that are difficult for the student—measured by a high loss or a low reward for example—the resulting synthetic data from the teacher is difficult as well, resulting in lower student accuracies on these generated samples than random selection.
At each iteration $t$ we use a given acquisition algorithm to select 11k samples $\bar{D}_{t}$ from $D_{t}$, before sending each to the teacher model to generate corresponding synthetic data $\hat{D_{t}}$. These data are appended to synthetic data from all previous iterations before reinitializing the student model and refitting its parameters with gradient descent.
@@ 5.1 Datasets
This section presents results on four distinct reasoning datasets in conjunction with four different models. GSM8k is a popular mathematics dataset comprised of school level maths problems (Cobbe et al., 2021), which we use with a Mistral-7B-Instruct-v0.3 student (Jiang et al., 2023a). Similarly, we include the more challenging Math1-3 dataset (Hendrycks et al., 2021), which is comprised of $5$ distinct levels of question difficulty—we use the easiest levels, $1$ to $3$, to finetune a Llama-3-8B-Instruct student (Dubey et al., 2024). We further experiment with the logical reasoning dataset ProntoQA (Saparov and He, 2023), composed of synthetically generated chain-of-thought style reasoning questions, with a Qwen1.5-7B-Chat student. Finally, we consider the Game of 24 dataset, which requires finding arithmetic operations to obtain $24$ given $4$ input integers. Here we use a Qwen2.5-7B-Instruct student (Qwen et al., 2025). Specifics are provided in Section D.1.
For all datasets except for Game of 24 we use prompt-based synthetic data generation with a GPT-4o teacher (prompts in  Section D.2). Instead, we use backward reasoning: if the answer is 13*8-10*8=24, for example, we can construct a new question by setting two integers to variables a*b-10*8=24 and solving to generate new questions (Jiang et al., 2024). We use a GPT-o3-mini teacher for backward reasoning, it qualitatively produces better question-response pairs than GPT-4o (Section D.2.4).
@@ 5.2 Finetuning Setup
Figure 2: SFT performance on 11k data points for various datasets and SLMs. We compare the effect of synthetic answer generation and synthetic question and answer generation to using the seed dataset, $D_{0}$ for SFT. 0-shot SLM and teacher performances are included for reference. All datasets use a GPT-4o teacher, for Game of 24 we use a GPT-o3-mini teacher. Using synthetic data, either as answers paired with real questions (syn. a) or both questions and answers (syn. q&a) improve performance past using the seed dataset alone (FT).
@@ 5.3 Algorithms
This paper considers a variety of selection algorithms. Prior work has shown that prioritizing “hard” samples accelerates learning (Section 3.1), which we also find to be the case for iterative synthetic data generation (Section C.6). This approach prioritizes high-uncertainty data, measured as the model’s loss on greedily decoded student generations, which we denote as “loss (high)” throughout this section. We also consider a low-reward selection algorithm (“rwd (low)”), also using the student’s own predictions, which scores generations using an external model. We use a Skywork-Reward-Llama-3.1-8B-v0.2 reward model which is the highest scoring 88b model on RewardBench (Lambert et al., 2025) at the time of writing.
We use Lion (Jiang et al., 2023b) as a baseline, which compares the student and teacher answer LLM-as-a-judge scores to classify each data point as either an easy or a hard before sampling equally from both sets. For completeness, we also consider a baseline that only samples from the hard set, denoted as LLM-as-a-judge (hard) (Jazbec et al., 2024). We use the same prompts for LLM-as-a-judge scoring as Jiang et al. (2023b).
We further consider prioritizing data with incorrect student answers, $s_{i}=\bm{1}\{\hat{y}_{i}\neq y\}$ as a proxy for prioritizing hard samples (Lee et al., 2024). In a similar spirit to Lion, we can instead sample evenly from correct and incorrect pools to maintain diversity in the seed data (Liu et al., 2024a). Since correct and incorrect scoring requires a verifier and ground-truth answers, we do not compare them to other scoring methods that do not use label information and instead place these supplementary results in Section C.1.
Figure 3: Student performance over successive synthetic data iterations with growing training sets. In all cases, selection based on uncertainty (loss) performs approximately as well as LLM-based scoring strategies (rwd and llm-judge), without requiring additional queries to an LLM. Further, for tasks that are out-of-distribution for the scoring model, like Game of 24, these mechanisms can perform even worse than random sampling. Horizontal lines in each inset plot denotes the proportion of data random sampling would require to achieve the same performance as the best active selection strategy in the corresponding experiment.